OpenStudy (anonymous):

Find the linear approximation of the function g(x)=(1+x)^(1/3) at a=0 and use it to approximate the numbers (0.95)^(1/3) and (1.1)^(1/3). Illustrate by graphing g and the tangent line.

OpenStudy (anonymous):

this is the same problem as "find the line tangent to the graph of \[g(x)=\sqrt[3]{x+1}\] at the point (0,1)

OpenStudy (anonymous):

you have the point, and to get the slope take the derivative and replace x by 0

OpenStudy (anonymous):

the derivative is \[g'(x)=\frac{1}{3\sqrt[3]{(x+1)^2}}\] and \[g'(0)=\frac{1}{3}\]

OpenStudy (anonymous):

so your line is \[y-1=\frac{1}{3}x

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

so your line is \[y-1=\frac{1}{3}x\] or \[y=\frac{1}{3}x+1\]

OpenStudy (anonymous):

if you want to approximate \[\sqrt[3]{.95}\]replace x by \[-.05\] this is some old school stuff now mainly of historical interest

OpenStudy (anonymous):

hope it is clear where the \[-.05\] came from i solved \[x+1=0.95\] for x

OpenStudy (anonymous):

what about the approximation of (1.1)^(1/3) and the graph?